# Variations of metric that preserve a Riemannian submersion and geometry of its fibers

### v1

@@ 4.2 Second variation
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal. Then for all $t\in(-\epsilon,\epsilon)$ the vector fields $\{P_{{\cal H}}^{t}\pi^{*}W_{n+1},\ldots,P_{{\cal H}}^{t}\pi^{*}W_{n+p}\}$ are $g_{t}$-orthonormal and $g_{t}$-horizontal, and
$\partial_{t}P_{{\cal H}}^{t}\pi^{*}W_{i}=-V_{i}$
@@
To prove the second one, we have for all $X\in\mathfrak{X}_{M}$:
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. If $g_{0}$ is a critical point of (29) ((((respectively, (30))))), with respect to variations preserving the Riemannian submersion $\pi$, then $g_{0}$ is a point of local minimum of (29) ((((respectively, (30))))), with respect to these variations.
@@
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a one-parameter family of Riemannian metrics on $M$ that preserve Riemannian submersion $\pi$ and let $x\in N$. By Remark 3.7, we can assume that (12) holds on $\pi^{-1}(U)$, for an open neighbourhood $U$ of $x$ in $N$, with $\{W_{n+1},\ldots,W_{n+p}\}$ that are $g_{N}$-orthonormal at $x$.
@@
$\partial_{t}h_{t}(X,Y)=\frac{1}{2}\sum\nolimits_{i=n+1}^{n+p}(\mathcal{L}_{V_{i}}g_{0})(X,Y)P_{{\cal H}}^{t}\pi^{*}W_{i}.$ (37)
@@
$\partial_{t}\|h_{t}\|^{2}=\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}}g_{0},h^{i}_{t}\rangle_{{\cal V}}$
and from (35) and (37) it follows that
$\partial_{t}h^{i}_{t}(X,Y)$ $=$ $\partial_{t}g_{t}(h_{t}(X,Y),P_{{\cal H}}^{t}\pi^{*}W_{i})=g_{t}(\partial_{t}h_{t}(X,Y),P_{{\cal H}}^{t}\pi^{*}W_{i})$
$=$ $\frac{1}{2}\sum\nolimits_{j=n+1}^{n+p}(\mathcal{L}_{V_{j}}g_{0})(X,Y)g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{j},P_{{\cal H}}^{t}\pi^{*}W_{i})$
$=$ $\frac{1}{2}(\mathcal{L}_{V_{i}}g_{0})(X,Y).$
@@
$\partial^{2}_{tt}\|h_{t}\|^{2}$ $=$ $\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{\partial_{t}V_{i}}g_{0},h^{i}_{t}\rangle_{{\cal V}}+\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}}g_{0},\partial_{t}h^{i}_{t}\rangle_{{\cal V}}$ (38)
$=$ $\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{\partial_{t}V_{i}}g_{0},h^{i}_{t}\rangle_{{\cal V}}+\frac{1}{2}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}}g_{0},\mathcal{L}_{V_{i}}g_{0}\rangle_{{\cal V}}.$
@@
$\partial_{t}H_{t}$ $=$ $\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{\ker\pi_{*}}V_{i})P_{{\cal H}}^{t}\pi^{*}W_{i},$
$\partial_{t}g_{t}(H_{t},H_{t})$ $=$ $2\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{\ker\pi_{*}}V_{i})g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{i},H_{t}),$
$\partial^{2}_{tt}g_{t}(H_{t},H_{t})$ $=$ $2\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{\ker\pi_{*}}\partial_{t}V_{i})g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{i},H_{t})$ (39)
$+2\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{\ker\pi_{*}}V_{i})^{2}.$
@@
$=\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{\partial_{t}V_{i}|_{t=0}}g_{0},h^{i}_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}+\frac{1}{2}\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}}g_{0},\mathcal{L}_{V_{i}}g_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}$
$=\frac{1}{2}\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}}g_{0},\mathcal{L}_{V_{i}}g_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}\geq 0.$ (40)
@@
$=2\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}g_{0}(\pi^{*}W_{i},H_{0})(\operatorname{div}^{0}_{\ker\pi_{*}}(\partial_{t}V_{i}|_{t=0}))\,\operatorname{vol}_{g_{0}}$
$\quad+2\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{\ker\pi_{*}}V_{i})^{2}\,\operatorname{vol}_{g_{0}}$
$=2\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{\ker\pi_{*}}V_{i})^{2}\,\operatorname{vol}_{g_{0}}\geq 0.$ (41)
@@
4.3 Sectional curvature
In this section we compute variations, with respect to metrics in $\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserving a Riemannian submersion $\pi$, of sectional curvatures in directions of horizontal lifts of given vector fields on $N$. Formulas for these variations are not tensorial and hence the critical points are only those metrics, for which the integrability tensor of the horizontal distribution vanishes on the lifts.
@@ 4.3 Sectional curvature
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal vector fields on $N$. Let $\sec_{M}(X,Y)$ be the sectional curvature of the plane field spanned by linearly independent $X,Y\in\mathfrak{X}_{M}$. Then
@@
$\quad-\frac{3}{2}g_{t}([V_{i},V_{j}]+[V_{i},V_{j}],P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}])$
$\quad-\frac{3}{2}g_{t}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]+[P_{{\cal H}}^{t}\pi^{*}W_{i},V_{j}],[V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]+[P_{{\cal H}}^{t}\pi^{*}W_{i},V_{j}]).$ (43)
@@
$-\frac{3}{4}g_{t}(P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i}m,P_{{\cal H}}^{t}\pi^{*}W_{j}],P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]).$ (48)
Using (4.3), we obtain (4.9) and (4.9) from (4.3) and (4.3). ∎
Setting $[W_{i},W_{j}]=0$, $\partial_{t}V_{i}=0$ and $[V_{i},V_{j}]=0$ for all $i,j\in\{n+1,\ldots,n+p\}$ in (4.9), we obtain the following.
@@
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal. If $[W_{i},W_{j}]=0$, $V_{i}$ do not depend on $t$ and $[V_{i},V_{j}]=0$ for all $i,j\in\{n+1,\ldots,n+p\}$, then $\partial^{2}_{tt}\sec_{M}(P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j})\leq 0$.

### v2

@@ 4.2 Second variation
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal. Then for all $t\in(-\epsilon,\epsilon)$ the vector fields $\{P_{{\cal H}}^{t}\pi^{*}W_{n+1},\ldots,P_{{\cal H}}^{t}\pi^{*}W_{n+p}\}$ are $g_{t}$-orthonormal and $g_{t}$-horizontal, and
$\partial_{t}P_{{\cal H}}^{t}\pi^{*}W_{i}=-V_{i}$ (38)
@@
To prove the last one, we have for all $X\in\mathfrak{X}_{M}$:
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. Let $g_{0}$ be a critical point of (30), ((respectively, (31))), with respect to variations preserving the Riemannian submersion $\pi$. Then for all such variations we have $\partial^{2}_{tt}J_{h}(g_{t})\geq 0$ ((respectively, $\partial^{2}_{tt}J_{H}(g_{t})\geq 0$ )).
@@
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a one-parameter family of Riemannian metrics on $M$ that preserve Riemannian submersion $\pi$ and let $x\in N$. By Remark 3.7, we can assume that (12) holds on $\pi^{-1}(U)$, for an open neighbourhood $U$ of $x$ in $N$, with $\{W_{n+1},\ldots,W_{n+p}\}$ that are $g_{N}$-orthonormal at $x$ and $t$-dependent $\{V_{n+1},\ldots,V_{n+p}\}$.
@@
$P_{{\cal H}}^{t}\partial_{t}h_{t}(X,Y)=\frac{1}{2}\sum\nolimits_{i=n+1}^{n+p}(\mathcal{L}_{V_{i}(t)}g_{0})(X,Y)P_{{\cal H}}^{t}\pi^{*}W_{i}.$ (41)
@@
$\partial_{t}\|h_{t}\|^{2}=\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}(t)}g_{0},h^{i}_{t}\rangle_{{\cal V}}$
and from (1), (39) and (41) it follows that
$\partial_{t}h^{i}_{t}(X,Y)$ $=$ $\partial_{t}g_{t}(h_{t}(X,Y),P_{{\cal H}}^{t}\pi^{*}W_{i})=g_{t}(\partial_{t}h_{t}(X,Y),P_{{\cal H}}^{t}\pi^{*}W_{i})$ (42)
$=$ $\frac{1}{2}\sum\nolimits_{j=n+1}^{n+p}(\mathcal{L}_{V_{j}(t)}g_{0})(X,Y)g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{j},P_{{\cal H}}^{t}\pi^{*}W_{i})$
$=$ $\frac{1}{2}(\mathcal{L}_{V_{i}(t)}g_{0})(X,Y).$
@@
$\partial^{2}_{tt}\|h_{t}\|^{2}$ $=$ $\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{\partial_{t}V_{i}}g_{0},h^{i}_{t}\rangle_{{\cal V}}+\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}(t)}g_{0},\partial_{t}h^{i}_{t}\rangle_{{\cal V}}$ (43)
$=$ $\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{\partial_{t}V_{i}}g_{0},h^{i}_{t}\rangle_{{\cal V}}+\frac{1}{2}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}(t)}g_{0},\mathcal{L}_{V_{i}(t)}g_{0}\rangle_{{\cal V}}.$
@@
$P_{{\cal H}}^{t}\partial_{t}H_{t}$ $=$ $\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{{\cal V}}V_{i}(t))P_{{\cal H}}^{t}\pi^{*}W_{i},$
$\partial_{t}g_{t}(H_{t},H_{t})$ $=$ $2\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{{\cal V}}V_{i}(t))g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{i},H_{t}),$
$\partial^{2}_{tt}g_{t}(H_{t},H_{t})$ $=$ $2\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{{\cal V}}\partial_{t}V_{i})g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{i},H_{t})$ (44)
$+2\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{{\cal V}}V_{i}(t))^{2}.$
@@
$=\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{\partial_{t}V_{i}|_{t=0}}g_{0},h^{i}_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}+\frac{1}{2}\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}(0)}g_{0},\mathcal{L}_{V_{i}(0)}g_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}$
$=\frac{1}{2}\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V_{i}(0)}g_{0},\mathcal{L}_{V_{i}(0)}g_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}\geq 0.$ (45)
@@
$=2\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}g_{0}(\pi^{*}W_{i},H_{0})(\operatorname{div}^{0}_{{\cal V}}(\partial_{t}V_{i}|_{t=0}))\,\operatorname{vol}_{g_{0}}$
$\quad+2\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{{\cal V}}V_{i}(0))^{2}\,\operatorname{vol}_{g_{0}}$
$=2\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}(\operatorname{div}^{0}_{{\cal V}}V_{i}(0))^{2}\,\operatorname{vol}_{g_{0}}\geq 0.$ (46)
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. Let $x\in N$ and let $g_{0}$ be a critical point of (36), with respect to variations preserving the Riemannian submersion $\pi$. Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation that preserves the Riemannian submersion $\pi$ and let $U$ be an open subset on $N$ such that $x\in U$ and (12) holds on $\pi^{-1}(U)$ for some $t$-dependent vector fields $V_{i}(t)$.
Let $q\geq 2$. We have $\frac{\partial^{r}}{\partial t^{r}}J_{h,x}(g_{t})|_{t=0}=0$ for all $1\leq r\leq q$ if and only if $\{V_{i}^{(s)}(0),i=n+1,\ldots,n+p\}$ are Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ for all $0\leq s\leq\lfloor\frac{q}{2}\rfloor-1$, where we use notation $V_{i}^{(s)}(0)=\frac{\partial^{s}}{\partial t^{s}}V_{i}|_{t=0}$ for all $s\geq 1$ and $V_{i}^{(0)}(0)=V_{i}(0)$.
Proof.
We have $\frac{\partial}{\partial t}J_{h,x}(g_{t})|_{t=0}=0$ by the assumption of $g_{0}$ being critical. For all $r\geq 2$, from differentiating (43) and using (42), it follows that $\frac{\partial^{r}}{\partial t^{r}}J_{h,x}(g_{t})|_{t=0}$ is a linear combination, with positive coefficients, of term
$\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V^{(r-1)}_{i}(0)}g_{0},h^{i}_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}},$ (47)
which vanishes by (34), and terms of the form
$\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V^{(\alpha)}_{i}(0)}g_{0},\mathcal{L}_{V^{(\beta)}_{i}(0)}g_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}},$ (48)
for all $\alpha,\beta\geq 0$ such that $\alpha+\beta=r-2$. In particular, in every term (48) we have either $\alpha\leq\frac{r}{2}-1$ or $\beta\leq\frac{r}{2}-1$, so if $\{V^{(s)}_{i}(0),i=n+1,\ldots,n+p\}$ are Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ for all $0\leq s\leq\lfloor\frac{r}{2}\rfloor-1$ then $\frac{\partial^{r}}{\partial t^{r}}J_{h,x}(g_{t})|_{t=0}=0$.
On the other hand, suppose that $\frac{\partial^{r}}{\partial t^{r}}J_{h,x}(g_{t})|_{t=0}=0$ for all $1\leq r\leq q$ for some $q\geq 2$. From (4.2) and $q\geq 2$, we obtain that $\{V_{i}(0),i=n+1,\ldots,n+p\}$ are Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$. Suppose that $\{V^{(\alpha)}_{i}(0),i=n+1,\ldots,n+p\}$ are Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$, for all $0\leq\alpha\leq s$, where $s\geq 0$. Then $\frac{\partial^{2s+2}}{\partial t^{2s+2}}J_{h,x}(g_{t})|_{t=0}$ is a linear combination, with positive coefficients, of term (47) with $r=2s+2$, which vanishes by (34); terms (48), where either $\alpha\leq s$ or $\beta\leq s$, which vanish by the assumption; and term
$\int_{{\cal F}_{x}}\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V^{(s+1)}_{i}(0)}g_{0},\mathcal{L}_{V^{(s+1)}_{i}(0)}g_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}.$
If $\frac{\partial^{2s+2}}{\partial t^{2s+2}}J_{h,x}(g_{t})|_{t=0}=0$, then the above term also vanishes, and hence $\{V^{(s+1)}_{i}(0),i=n+1,\ldots,n+p\}$ are Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$. By induction on $s$, we obtain that $\{V^{(\alpha)}_{i}(0),i=n+1,\ldots,n+p\}$ are Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ for all $\alpha\leq s$, as long as $2s+2\leq q$, i.e., $s\leq\lfloor\frac{q}{2}\rfloor-1$. ∎
Remark 4.10.
A theorem analogous to Theorem 4.9 can be proved for functional (37), with $J_{h,x}$ replaced by $J_{H,x}$ and words “Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$” replaced by “divergence-free vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$”.
5 Sectional curvature
5.1 Horizontal curvature
In this section we compute variations, with respect to metrics in $\operatorname{Riem}(M,{\cal V},g_{0})$ preserving a Riemannian submersion $\pi$, of sectional curvatures in directions of horizontal lifts of given vector fields on $N$. Formulas for these variations are not tensorial and hence the critical points are only those metrics, for which the integrability tensor of the horizontal distribution vanishes on the lifts.
Theorem 5.1.
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal vector fields on $N$. Let $\sec_{M}(X,Y)$ be the sectional curvature of the plane field spanned by linearly independent $X,Y\in\mathfrak{X}_{M}$. Then
@@ 5.1 Horizontal curvature
$\quad-3g_{t}([V_{i},V_{j}],P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}])$
$\quad-\frac{3}{2}g_{t}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]+[P_{{\cal H}}^{t}\pi^{*}W_{i},V_{j}],[V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]+[P_{{\cal H}}^{t}\pi^{*}W_{i},V_{j}]).$
@@
$-\frac{3}{4}g_{t}(P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]).$ (55)
Using (5.1), we obtain (5.1) and (50) from (5.1) and (5.1). ∎
Setting $W_{i},W_{j}]=0$, $\partial_{t}V_{i}=0$ and $V_{i},V_{j}]=0$ for all $i,j\in\{n+1,\ldots,n+p\}$ in (50), we obtain the following.
@@
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal. If $W_{i},W_{j}]=0$, $V_{i}$ do not depend on $t$ and $V_{i},V_{j}]=0$ for all $i,j\in\{n+1,\ldots,n+p\}$, then $\partial^{2}_{tt}\sec_{M}(P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j})\leq 0$.
@@
Remark 5.4.
We note that for a local $g_{N}$-orthonormal frame $\{W_{n+1},\ldots W_{n+p}\}$ on $N$ we can define on $(M,g_{t})$ the scalar $g_{t}$-horizontal curvature $K_{{\cal H}(t)}=\sum\nolimits_{i,j=n+1}^{n+p}\sec(P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j})$. By (5.1), we have $K_{{\cal H}(t)}=K_{N}-\frac{3}{4}\|{\tilde{T}}_{t}\|^{2}$, where $K_{N}$ is the scalar curvature of $(N,g_{N})$ and $\|{\tilde{T}}_{t}\|^{2}=\sum\nolimits_{i=n+1}^{n+p}g_{t}(P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}])$. Thus, varying $K_{{\cal H}(t)}$ is equivalent to varying $\|{\tilde{T}}_{t}\|^{2}$. Examples of critical points of functional $g_{t}\mapsto\int_{M}\|{\tilde{T}}_{t}\|^{2}\,\operatorname{vol}_{g_{t}}$ on a domain of a Riemannian submersion $\pi:(M,g)\rightarrow(N,g_{N})$ are metrics, for which $(M,g)$ is a $3$-Sasakian (which was proved in [16]), or a $f$-$K$-contact manifold (which can be proved by similar computations as those in [18]), in each case with vertical characteristic vector fields.
5.2 Vertizontal curvature of totally geodesic fibers
The curvatures $\sec_{M}(X,U)$ for horizontal $X$ and vertical $U$ are called vertizontal, and a Riemannian submersion with totally geodesic fibers and all vertizontal curvatures positive is called fat [20]. In this section we compute variations of vertizontal curvature preserving a Riemannian submersion with totally geodesic fibers and show that on certain manifolds (e.g., spheres $S^{4m+3}$ with the Hopf fibration by $S^{3}$) we can make this variation non-constant on a fiber.